In the 1970's, when lead was was widely used in "ethyl" gasoline, human blood levels were such that the body of an average American contained 0.22 PPM lead. The danger for lead poisoning is 0.80 ppm. (a) What percent of the average person is lead? (b) How much lead would be in an average 79 kg person? (c) How much more lead would the average person need to accumulate to reach the danger level?

A- ___ x 10 ^ %

B. ____ g

C. ____ g
Answer #1

so 0.22 ppm = 2.2* 10^-4 %

lead in 79 kg person = 2.2* 10^-4 * 79 / 100 = 0.1738 gram

for danger levels more lead required = 8* 10^-4 - 2.2 * 10^4 % = 5.8* 10^-4 % = 5.8* 10^-6 * 79 = 0.4582g